OpenStudy (anonymous):

c=−30x2+1140x+9ˏ450 |

OpenStudy (anonymous):

mhm

OpenStudy (jdoe0001):

\(\bf c=−30x^2+1140x+9ˏ450\implies c = 30(x^2-38x+315)\)

OpenStudy (jdoe0001):

hmm forgot the minus \(\bf c=-30x^2+1140x+9ˏ450\implies c = -30(x^2-38x+315)\) so... there, to find "x", set c = 0 and factor the quadratic